one leg of a triangle is 1 centimeter shorter than the other leg and the hypotenuse is 2 centimeters longer than the longer leg. What are the lengths of the sides of the triangle?
Answer by mananth(12270)

let one leg be x
other leg = x-1
hypotenuse = x+2
..
x^2+(x-1)^2=(x+2)^2
x^2+x^2-2x+1= x^2 +4x+4
x^2-6x-3=0
solve for x using quadratic formula
discriminant = 48
x1=((6+sqrt(48))/2
x1= 6.46cm one leg
other leg will be 5.46cm
hypotenuse = 7.46cm
...
m.ananth@hotmail.ca